OpenStudy (anonymous):

Find all real solutions of the equation. (Enter your answers as a comma-separated list. If there is no real solution, enter NO REAL SOLUTION.) x5 − 625x = 0

OpenStudy (anonymous):

(0,5)

OpenStudy (anonymous):

That is x=0 or x=5

OpenStudy (anonymous):

what about this one : Find all real solutions of the equation. (Enter your answers as a comma-separated list. If there is no real solution, enter NO REAL SOLUTION.) x3 − x2 + 5x − 9 = x2 + 1

OpenStudy (anonymous):

bring everything over to one side, group in pairs.

OpenStudy (anonymous):

x^3 -2x^2 +5x -10 =0 x^2 ( x-2) +5(x-2) =0 (x^2 +5)(x-2) =0

OpenStudy (anonymous):

so the only answer is 2

OpenStudy (anonymous):

?

OpenStudy (anonymous):

yes